# Exploring dynamical quantum phase transition from pure states to mixed states through extended Su-Schrieffer-Heeger models

### v2

@@ IV DQPT for mixed states in SSH-3 model
In this section, we investigate the effect of the finite temperature on the DQPT, in which the generalized Loschmidt amplitude for mixed states is calculated. Fig. 7 shows the rate functions in the SSH-3 model for initial states at different temperatures. It can be observed that at very low temperatures, i.e. $T=0.1$ and $T=0.2$, the rate function almost coincides with the result of that starting from the initial pure states $|u_{k1}^{i}\rangle$. This indicates that the behavior of the DQPT is not affected at low temperatures. As the temperature increases, the critical time will gradually deviate from the critical time at pure states and lower temperatures. Especially, at $T=0.5$, the numerical values of the rate function change but the critical time still equals that at lower temperatures. However, when the temperature is very high, the nature of DQPTs will undergo significant changes, to the extent that we observe a large number of critical times emerging in the system through the first derivative $d\lambda/dt$ of the rate function. In addition, we also observe the same behavior of DQPTs from the Pancharatnam geometric phase (see Fig. 7). In particular, at high temperature, i.e. $T=2.0$, the Pancharatnam geometric phase can be seen to exhibit many dynamical vortices emerging in the ($k-t$) plane.
@@
The behavior of the DQPT for mixed states at finite temperatures can be explained based on the properties of that for pure states. At lower temperatures, the probabilities $f_{k\mu}$ of the thermal states satisfy
@@
thus the generalized Loschmidt amplitude is approximately equal to the case starting from the lowest energy band in the pure state. As the temperature increases, the probabilities $f_{k\mu}$ of the thermal states change to
@@
so that the generalized Loschmidt amplitude is approximately equal to the cases only mixing the quench from pure states $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$. According to Fig. 3, we know that the cases of quenches from initial states $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$ have the common critical wave vectors $k_{c1}$ and $-k_{c1}$. In this case, the critical times for mixed states still do not change or just deviate slightly from that at lower temperatures. However, at high temperatures, due to the contribution of that from the highest band $|u_{k3}^{i}\rangle$ no longer being negligible, and considering the cases of quenches from the initial pure states $|u_{k2}^{i}\rangle$ and $|u_{k3}^{i}\rangle$ have another common critical time $k_{c2}(-k_{c2})$, which differs from $k_{c1}$, the generalized Loschmidt amplitude $\mathbb{G}(t)$ will not have the zeros at $k_{c1}$ and $k_{c2}$. Consequently, the behavior of the DQPT at high temperatures is significantly changed.
V DQPTs in SSH-4 model
Figure 8: The energy spectra as functions of $t_{4}$ in the SSH-4 models for (a) with $t_{1}=t_{2}=t_{3}=1$, and for (b) with $t_{1}=t_{3}=1$ and $t_{2}=2$.
@@ V DQPTs in SSH-4 model
Figure 9: The rate functions in the SSH-4 model with the energy spectrum configuration corresponding to Fig. 8 (a) for the quenches from different initial states. The quench path in (a) is from $t_{4}^{i}=0.2$ to $t_{4}^{f}=0.8$, and in (b) is from $t_{4}^{i}=0.2$ to $t_{4}^{f}=1.2$.
As another typical example, we discuss the dynamic behavior in the SSH-4 model in this section. Unlike the SSH-3 model, the Bloch Hamiltonian of the SSH-4 model satisfies the chiral symmetry, leading to the energy spectrum of the SSH-4 being symmetric about the transverse axis ($\varepsilon_{k}=0$) for a specific set of parameters, i.e.
$\varepsilon_{k1}=-\varepsilon_{k4},\quad\varepsilon_{k2}=-\varepsilon_{k3}.$ (25)
By calculating the energy spectra under different parameters, we have identified three typical energy spectrum configurations in the SSH-4 model, one of which resembles the energy spectra of the period-two Kitaev chain, as discussed in the Ref. Cao et al. (2024b). In this article, we will discuss two other cases different from the period-two Kitaev chains, with the shapes of their energy spectra shown in Fig. 8.
Figure 10: The rate functions in the SSH-4 model with the energy spectrum configuration corresponding to Fig. 8 (b) for the quenches from different initial states. The quench path in (a) crossing the first critical point $t_{4c}=0.5$ is from $t_{4}^{i}=0.2$ to $t_{4}^{f}=0.8$, and in (b) also crossing the second critical point $t_{4c}=2$ is from $t_{4}^{i}=0.2$ to $t_{4}^{f}=2.2$.‘The first deviation $d\lambda/dt$ of the rate function for the quench from the initial state $|u_{k1}^{i}\rangle$ is shown to see the critical time more clearly.
From Fig. 8 (a), it is can be seen the system with parameters $t_{1}=t_{2}=t_{3}$ has a gap-closing point at $t_{4c}=1$. In contrast, as shown in Fig. 8 (b), the system with parameters $t_{1}=t_{3}=1$ and $t_{2}=2$ has two gap-closing point at $t_{4c}=0.5$ and $t_{4c}=2$. These critical points that collapse the energy bands will result in the topological transitions in the system, which are characterized by the change in the number of edge modes Lee et al. (2022). It is evident that based on previous experience, we can speculate that in the SSH-4 model, the behavior of the DQPT after a quench will depend on whether the quench passes through these critical points.
@@
Figs. 9 and 10 present the rate functions obtained from quenches starting from different initial states in the SSH-4 model with two different energy spectrum configurations, respectively. It is worth noting that due to the chiral symmetry resulting in a symmetric relationship between energy spectra, as shown in Eq. (25), the rate functions for quenches from the initial bands $|u_{k3}^{i}\rangle$ and $|u_{k4}^{i}\rangle$ are identical to those from $|u_{k2}^{i}\rangle$ and $|u_{k1}^{i}\rangle$. Therefore, it is sufficient to show the results of the quenches from $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$ in the SSH-4 model. By comparing these results, we can verify our previous speculation that the system will only undergo a DQPT when the quench passes through the critical points. Interestingly, since the critical point at $t_{4c}=0.5$ corresponds to the closing of the energy gaps between bands $\varepsilon_{k2}$ and $\varepsilon_{k3}$, the quench from the initial state $|u_{k1}^{i}\rangle$ does not induce a DQPT even if it only crosses the critical point $t_{4c}=0.5$ [see Fig. 10 (a)].
@@
Some details in the results are also worth noting. In the SSH-3 model, we observe that the critical times obtained from quenches starting from different initial states are the same. However, in the SSH-4 model, the situation is different. From Figs. 9 (b) and 10 (b), it can be seen that the critical times of the rate functions for the quenches starting from $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$ are not equal. This difference is also reflected in the deviation of the critical wave vectors [see Fig. 1 (a) and (b)]. While these details may seem less important in the context of pure-state DQPTs, they will significantly alter the behavior of mixed-state DQPTs.
@@
Figure 12: The rate functions in the SSH-4 model for mixed states at different temperatures. The results in (a) correspond to the quench path of Fig. 9 (b), and in (b) correspond to that of Fig. 10 (b). The red-dot line denotes the rate function for the initial state $|u_{k1}^{i}\rangle$.
To investigate the influence of finite temperature on DQPTs, we present in Fig. 12 the rate functions obtained from mixed states at different initial temperatures. It can be observed that when the temperature is relatively low, the rate function aligns with the results from pure states. However, at higher temperatures, the rate function becomes smooth, no longer exhibiting non-analytic critical points. This indicates that the DQPT in mixed states disappears at high temperatures. Similar to the discussion in the SSH-3 model, we can also understand the behavior of dynamic quantum phase transitions in mixed states at finite temperatures based on the behavior of pure-state DQPTs. When the temperature is low, the probabilities of different thermodynamic states satisfy $f_{k1}\gg f_{k2}>f_{k3}>f_{k4}$, thus the generalized Loschmidt amplitude of the system is approximately equal to the Loschmidt amplitude from quenching starting from the lowest energy band in the pure state. Therefore, at low temperatures, the behavior of DQPTs is the same as that of pure states. However, at high temperatures, due to the contribution of higher energy bands that cannot be neglected and the different critical wave vectors in quenching processes starting from various bands, the generalized Loschmidt amplitude does not exhibit Fisher zeros, leading to the disappearance of DQPTs in mixed states.
VI conclusion
@@
In this paper, we explore DQPTs for both pure and mixed states within the generalized SSH model, focusing on two representative configurations: the SSH-3 and SSH-4 models, which exhibit different symmetries. The SSH-3 model does not possess chiral symmetry; instead, it features a chiral-like symmetry known as point symmetry, which gives rise to robust, localized edge states linked to the system’s topological properties. Our findings indicate that DQPTs for the pure state occur after a quench that crosses the topological transition, even when the energy band gap is open. In the case of mixed states, the DQPT behavior remains unaffected at low temperatures. However, significant changes in DQPT characteristics arise at high temperatures, leading to the emergence of multiple critical times within the system. Conversely, the SSH-4 model exhibits chiral symmetry and allows us to examine two distinct energy spectrum configurations. We find that the occurrence of DQPTs for the pure state requires the quench to start from an initial state without a band gap and to cross the critical point of the topological transition. For mixed states, DQPTs are absent at high temperatures.
@@ Acknowledgements.
K. Cao thanks Mrs. H. Zhou and prof. P. Tong for the valuable suggestion. The work is supported by the National Natural Science Foundation of China (Grant No. 11875047).

### v3

@@ III.3 DQPT for mixed states in SSH-3 model
In this section, we investigate the influence of the finite temperature on the DQPT in the SSH-3 model, in which the dynamics is described by the generalized Loschmidt amplitude for a quench from mixed initial states. First, it is necessary to clarify that since we find similar behavior of mixed-state DQPTs in the SSH-3 model both with and without inversion symmetry, the following results will use the system with inversion symmetry as an example to illustrate the nature of DQPTs at finite temperature.
Fig. 6 shows the rate functions in the SSH-3 model for initial states at different temperatures ($T=1/\beta$). It can be observed that at very low temperatures, i.e. $T=0.1$ and $T=0.2$, the rate function almost coincides with the result of that starting from the initial pure states $|u_{k1}^{i}\rangle$. This indicates that the behavior of the DQPT is not affected at low temperatures. As the temperature increases, the critical time will gradually deviate from the critical time at pure states and lower temperatures. Especially, at $T=0.5$, the numerical values of the rate function change but the critical time still equals that at lower temperatures. However, when the temperature is very high, the nature of DQPTs will undergo significant changes, to the extent that we observe a large number of critical times emerging in the system through the first derivative $d\lambda/dt$ of the rate function. In addition, the results of DQPTs are also confirmed through the Pancharatnam geometric phase (see Fig. 7). Especially, at high temperature $T=2.0$, the Pancharatnam geometric phase exhibits many dynamical vortices emerging in the ($k-t$) plane.
@@
The behavior of the DQPT for mixed states at finite temperatures can be explained based on the properties of that from pure states. At lower temperatures, the probabilities $f_{k\mu}$ of the thermal states satisfy
@@
thus the generalized Loschmidt amplitude is approximately equal to the case starting from the lowest energy band in the pure state [see results of $T=0.1,0.2$ in Fig. 7 (a)]. As the temperature increases, the probabilities $f_{k\mu}$ of the thermal states shift to
@@
so that the generalized Loschmidt amplitude is approximately equal to the cases only mixing the quench from pure states $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$. According to Fig. 2, we know that the cases of quenches from initial states $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$ have the common critical wave vectors $k_{c1}$ and $-k_{c1}$. In this case, the critical times for mixed states still do not change or just deviate slightly from that at lower temperatures [see results of $T=0.5,1.0$ in Fig. 7 (a)]. However, at high temperatures, due to the contribution of that from the highest band $|u_{k3}^{i}\rangle$ no longer being negligible. Considering the cases of quenches from the initial pure states $|u_{k2}^{i}\rangle$ and $|u_{k3}^{i}\rangle$ have another common critical time $k_{c2}(-k_{c2})$, distinct from $k_{c1}$, the behavior of the DQPT at high temperatures is significantly changed [see results of $T=2.0$ in Fig. 7 (b)].
IV DQPTs in SSH-4 model
Figure 8: Energy spectra as a function of the intercell coupling $t_{4}$ in the SSH-4 models, for (a) with the intracell couplings $t_{1}=t_{2}=t_{3}=1$, for (b) with $t_{1}=t_{3}=1$ and $t_{2}=0.5$, for (c) with $t_{1}=t_{3}=2$ and $t_{2}=1$, and for (d) with $t_{1}=1,t_{2}=2,t_{3}=3$. The solid lines denotes the bulk states, the dotted lines represent the edge states. The nonzero edge modes are given by $\varepsilon_{k}=\pm\sqrt{t_{1}^{2}+t_{2}^{2}}$. For convenience, we label the phases using Roman letters.
[…]
By calculating the energy spectra under different configurations of couplings ($t_{1},t_{2},t_{3},t_{4}$), we identify three typical energy spectrum configurations, characterized by the different locations of edge states, in the SSH-4 model. If all the intracell couplings are equal, e.g. $t_{1}=t_{2}=t_{3}=1$, the Hamiltonian exhibits the band-gap closing between two neighboring energy spectra at the critical point $t_{4c}=1$ [see Fig. 8 (a)]. Three edge modes (one zero edge mode and two nonzero edge modes $\varepsilon_{k}=\pm\sqrt{t_{1}^{2}+t_{2}^{2}}$) are present in the phase with $t_{4}>1$. If the intracell couplings satisfy $t_{1}=t_{3}\neq t_{2}$, the Hamiltonian has two critical points $t_{4c}$, at which the band gap closes [see Figs. 8 (b) and (c)]. The two critical points are obtained by
$t_{4c}=\frac{t_{1}t_{3}}{t_{2}},\quad\frac{t_{2}t_{3}}{t_{1}},$ (25)
respectively. In this case, the locations of edge states are determined by $t_{2}$: if $t_{2}<t_{1}(t_{3})$, the Hamiltonian has two nonzero edge modes for $\frac{t_{2}t_{3}}{t_{1}}<t_{4}<\frac{t_{1}t_{3}}{t_{2}}$, and three edge modes (one zero edge mode and two nonzero edge modes) for $t_{4}>\frac{t_{1}t_{3}}{t_{2}}$ [see Fig. 8 (b)]; if $t_{2}>t_{1}(t_{3})$, the Hamiltonian has one zero edge mode for $\frac{t_{1}t_{3}}{t_{2}}<t_{4}<\frac{t_{2}t_{3}}{t_{1}}$, and three edge modes (one zero edge mode and two nonzero edge modes) for $t_{4}>\frac{t_{2}t_{3}}{t_{1}}$ [see Fig. 8 (c)]. Finally, if the intracell couplings satisfy $t_{1}\neq t_{3}$, the Hamiltonian has only one critical point at
$t_{4c}=\frac{t_{1}t_{3}}{t_{2}},$ (26)
and one zero edge mode is present for $t_{4}>\frac{t_{1}t_{3}}{t_{2}}$.
IV.1 DQPTs for pure states in SSH-4 model
@@
Figure 9: Rate functions in the SSH-4 model with different energy spectra configurations for the quench from the initial pure states $|u_{k1}^{i}\rangle(|u_{k4}^{i}\rangle)$.
Figure 10: Rate functions in the SSH-4 model with different energy spectra configurations for the quench from the initial pure states $|u_{k2}^{i}\rangle(|u_{k3}^{i}\rangle)$.
Now we investigate the DQPTs in the SSH-4 model from the pure states. First, it should be noted that, similar to the Kitaev chain which possesses particle-hole symmetry Cao et al. (2024b), the SSH-4 model exhibits chiral symmetry, resulting in the system’s energy spectrum satisfying $\varepsilon_{k1}=-\varepsilon_{k4}$ and $\varepsilon_{k2}=-\varepsilon_{k3}$. Moreover, along the same quench pathway, the rate functions obtained from quantum quenches starting from $|u_{k1}^{i}\rangle$ and $|u_{k4}^{i}\rangle$, as well as from $|u_{k2}^{i}\rangle$ and $|u_{k3}^{i}\rangle$, are exactly the same. Therefore, here we only need to consider the results starting from $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$.
Figs. 9 show the typical results of rate functions for various types of quenches starting from the initial pure states $|u_{k1}^{i}\rangle$ in the SSH-4 model, under different energy spectrum configurations. We observe that quenches from phase I to phase II, phase III to phase IV, and phase VII to phase VIII exhibit cusp-like singularities in the rate functions, indicating the presence of dynamical quantum phase transitions (DQPTs). In contrast, other quenches crossing critical points show smooth rate functions without DQPT signatures. Notably, the quenches accompanied by DQPTs correspond to Hamiltonians where the energy gap closes between bands $\varepsilon_{k1}$ and $\varepsilon_{k2}$. These results suggest two necessary conditions for the occurrence of DQPTs: (1) the energy band of the initial state must have a gapless point with other bands, and (2) the quench must cross this gapless point. To verify this, Fig. 10 shows the rate functions for the same quenches starting from the initial states $|u_{k2}^{i}\rangle$ in Fig. 10. As illustrated in Fig. 8, critical points always exist where the energy gap between $\varepsilon_{k2}$ and other bands ($\varepsilon_{k1}$ or $\varepsilon_{k3}$) closes. Consequently, we observe that the quenches from $|u_{k2}^{i}\rangle$ exhibit cusp-like singularities, confirming the essential role of gap closing in DQPT emergence.
@@
Some details in the results are also worth noting. In the SSH-3 model, we observe that the critical times obtained from quenches starting from different initial states are the same. However, in the SSH-4 model, the situation is different. From Figs. 9 and 10, it can be seen that the critical times of the rate functions for the quenches starting from $|u_{k1}^{i}\rangle$ and $|u_{k2}^{i}\rangle$ are not equal ($t_{c}=6.15$ for $|u_{k1}^{i}\rangle$ and $t_{c}=5.82$ for $|u_{k2}^{i}\rangle$). This difference is also reflected in the deviation of the critical wave vectors [see Fig. 1 (a) and (b)]. While these details may seem less important in the context of pure-state DQPTs, they will significantly alter the behavior of mixed-state DQPTs.
IV.2 DQPTs for mixed states in SSH-4 model
@@
Figure 12: Rate functions in the SSH-4 model for mixed states at different temperatures. The red-dot line denotes the rate function for the initial state $|u_{k1}^{i}\rangle$ at zero temperature.
Figure 13: Rate functions in the SSH-4 models for various types of quenches at $T=0.1$.
We now turn to investigating the influence of finite temperature on DQPTs, where the quench begins from mixed initial states. Broadly, two scenarios arise for DQPTs in mixed states: one in which a quench starting from the pure state $|u_{k1}^{i}\rangle$ induces a DQPT, and another where it does not.
Figs. 12 (a) and (b) show the rate functions at various temperatures for quenches from phase I to phase II and from phase III to phase IV, respectively. For comparison, the rate function results for the corresponding initial pure state $|u_{k1}^{i}\rangle$ are also depicted. At relatively low temperatures, the rate functions closely follow those of the pure states. However, as temperature increases, the rate functions become smooth and no longer display non-analytic critical points, indicating the suppression of DQPTs in mixed states at elevated temperatures.
Similar to observations in the SSH-3 model, the finite-temperature behavior of DQPTs in SSH-4 model for mixed states can be understood by considering the pure-state DQPTs. At low temperatures, the thermodynamic occupation probabilities satisfy
$f_{k1}\gg f_{k2}>f_{k3}>f_{k4},$ (27)
thus the generalized Loschmidt amplitude of the system is approximately equal to the Loschmidt amplitude from quenching starting from the lowest energy band ($|u_{k1}^{i}\rangle$) closely approximating the pure-state Loschmidt amplitude. Consequently, DQPT behavior at low temperatures mirrors that of the pure state. At higher temperatures, however, the increased population of higher energy bands—each characterized by distinct critical momenta in the quench dynamics [see Fig. 11 (a) and (b)]—results in a generalized Loschmidt amplitude free of Fisher zeros, thereby eliminating DQPTs in the mixed state.
Conversely, if a quench starting from the pure state $|u_{k1}^{i}\rangle$ does not induce a DQPT, then DQPTs remain absent in the corresponding mixed-state quenches at all temperatures, including low temperature regimes [see Fig. 13].
V conclusion
@@
In this paper, we investigate DQPTs for both pure and mixed states within generalized SSH models, focusing on two representative cases: the SSH-3 and SSH-4 models, which differ in their symmetry properties. The SSH-3 model lacks chiral symmetry but possesses a chiral-like, point symmetry that protects robust, localized edge states associated with the system’s topology. Our results show that, for pure states, DQPTs occur following quenches that cross the topological phase transition—even when the energy band gap remains open. For mixed states, DQPT behavior at low temperatures closely follows that of pure states, while at higher temperatures significant modifications arise, including the appearance of multiple critical times. In contrast, the SSH-4 model, which exhibits chiral symmetry, allows exploration of four distinct energy spectrum configurations. We find that pure-state DQPTs require the quench to initiate from an initial state with a gapless band and to cross the critical point of the topological transition. At finite temperature, when a quench from the lowest energy band induces a DQPT, mixed-state DQPTs persist at low temperature but vanish at elevated temperatures. Conversely, if a quench starting from the pure state $|u_{k1}^{i}\rangle$ fails to induce a DQPT, no DQPTs emerge for mixed states at any temperature.
@@ Acknowledgements.
K. Cao thanks Mrs. H. Zhou and Prof. P. Tong for the valuable suggestions. The work is supported by the National Natural Science Foundation of China (Grant No. 11875047).